# Yaohui Li2, Jun Lan1, Huijia Zhu1, Jianfu Zhang3, Weiqiang Wang1,Huaxiong Li2 1Ant Group

### v2

@@ 3.3 Video collection details of GenVideo
Fake video collection: The guiding principle for collecting fake videos is to ensure maximal diversity in content and generators. We prioritize generating additional fake videos using the most recent generators due to their superior quality. To collect diverse fake videos from different resources as training samples, we have established a video generation pipeline for text-to-video generation and image-to-video generation. This pipeline facilitates the production of videos using popular generative mechanisms, including diffusion-based models (Xing et al., 2023; Zhang et al., 2023b; Guo et al., 2023a; Wang et al., 2023c; Blattmann et al., 2023b; Chen et al., 2024; Xing et al., 2023; Zhang et al., 2023b; Chen et al., 2023c), transformer-based methods (Ma et al., 2024b; Ope, 2024) and service-based method Pika website (pik, 2022). We selected common categories such as humans, animals, and plants as foreground keywords, and typical scenes like “in the park” or “on the lawn” as background keywords. Leveraging a large language model (Le Scao et al., 2023), these foreground and background keywords are expanded into comprehensive textual prompts to guide text-to-video generation. For image-to-video generation, we employed various text-to-image models, including different versions of Stable Diffusion (SD (Rombach et al., 2022), SDXL (Podell et al., 2023)). Using the enriched textual prompts generated by the large language model, we created corresponding images, which were subsequently used as input for image-to-video models to generate the final videos.
@@ 4.2 AI-generated video detection with DeMamba module
Figure 2: The overall framework of our Detail Mamba (DeMamba).
@@ 5.1 Implementation details
Task Model Batchsize LR Samples Epochs Optimizer Scheduler Loss
per epoch
F3Net 1024 1e-5 $1024\times 2000$ 30 AdamW [20, 25], lr * 0.1 BCE
NPR 1024 1e-5 $1024\times 2000$ 30 [20, 25], lr * 0.1
STIL 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
VideoMAE-B 128 1e-5 $128\times 4000$ 30 [20, 25], lr * 0.1
many- CLIP-B-LP 1024 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
to- CLIP-B-FT 1024 1e-6 $1024\times 2000$ 10 -
many XCLIP-B-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
generalization XCLIP-B-FT 128 1e-6 $128\times 4000$ 10 -
DeMamba-CLIP-LP 128 1e-6 $1024\times 2000$ 30 [20, 25], lr * 0.1
DeMamba-CLIP-FT 128 1e-6 $1024\times 2000$ 10 -
DeMamba-XCLIP-LP 128 1e-6 $128\times 4000$ 30 [20, 25], lr * 0.1
DeMamba-XCLIP-FT 128 1e-6 $128\times 4000$ 10 -
one- STIL 128 1e-5 $128\times 400$ 10 AdamW - BCE
to- NPR 1024 1e-5 $1024\times 2000$ 10 -
many XCLIP-B-FT 1024 1e-6 $128\times 400$ 10 -
generalization DeMamba-XCLIP-FT 128 1e-6 $128\times 400$ 10 -
Table 3: Comparisons to the SOTAs in mean accuracy (ACC) and average precision (AP) on the many-to-many generalization task.
Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real Avg.
@@
F3Net Image ACC 83.93 99.71 98.62 77.57 57.00 36.57 99.52 99.71 89.43 76.78 99.14 83.45
AP 68.27 99.89 99.67 89.35 85.24 63.17 99.58 99.89 93.80 88.41 - 88.73
NPR Image ACC 91.07 99.57 99.49 24.29 89.64 57.71 97.12 99.86 94.29 87.80 97.46 85.30
AP 67.17 99.14 99.20 22.76 93.91 61.76 96.33 99.72 94.15 90.40 - 82.45
STIL Video ACC 78.57 98.14 98.04 76.00 61.79 53.29 99.36 97.36 94.57 65.01 98.72 83.71
AP 57.21 99.08 99.32 86.19 82.24 70.43 99.25 98.96 97.18 81.32 - 87.12
VideoMAE-B Video ACC 67.86 96.00 98.41 96.14 77.14 80.43 97.44 96.93 96.29 68.36 99.71 88.61
AP 66.49 98.85 99.77 99.27 96.55 95.31 99.49 99.69 99.27 90.74 - 94.54
CLIP-B-PT Image ACC 85.71 82.43 90.36 71.00 79.29 75.43 89.62 86.29 82.14 75.16 57.22 79.67
AP 6.78 43.56 70.88 29.97 52.97 35.36 55.52 66.03 44.23 42.99 - 44.83
DeMamba Video ACC 58.93 96.43 93.12 68.00 69.36 69.00 89.14 91.86 96.14 56.59 98.06 $80.60_{{\color[rgb]{0.453125,0,0}+0.93}}$
-CLIP-PT AP 25.87 95.14 96.23 73.43 83.31 75.49 90.17 95.06 95.05 69.95 - $79.97_{{\color[rgb]{0.453125,0,0}+35.14}}$
CLIP-B-FT Image ACC 94.64 99.86 91.38 77.29 88.14 86.00 99.68 99.79 84.29 84.67 97.38 91.19
AP 80.67 99.67 95.24 82.20 93.48 88.62 99.55 99.79 86.93 89.08 - 91.52
DeMamba Video ACC 95.71 100.00 98.70 69.14 92.43 93.29 100.00 100.00 83.57 82.94 99.44 $92.29_{{\color[rgb]{0.453125,0,0}+1.10}}$
CLIP-B-FT AP 85.50 100.00 99.59 76.15 96.78 96.99 99.97 100.00 89.80 89.72 - $93.45_{{\color[rgb]{0.453125,0,0}+1.93}}$
XCLIP-B-PT Video ACC 81.34 82.15 83.35 80.98 81.82 81.55 82.14 82.98 81.93 81.10 81.37 81.88
AP 16.39 72.16 87.77 39.86 65.57 54.26 75.23 84.80 61.60 55.28 - 61.29
DeMamba Video ACC 66.07 95.86 94.64 77.86 75.36 80.29 90.89 92.50 96.00 66.41 95.12 $84.64_{{\color[rgb]{0.453125,0,0}+2.76}}$
-XCLIP-PT AP 18.26 93.50 94.72 69.94 78.08 71.50 83.95 92.23 93.54 68.10 - $76.38_{{\color[rgb]{0.453125,0,0}+15.09}}$
XCLIP-B-FT Video ACC 82.14 99.57 93.62 61.29 79.36 69.71 97.92 99.79 77.14 83.59 98.14 85.66
AP 64.42 99.73 96.78 70.98 90.35 77.28 97.34 99.84 82.01 88.97 - 86.77
DeMamba Video ACC 98.21 100.00 99.86 65.43 94.86 98.86 100.00 100.00 92.86 89.09 99.42 $\textbf{94.42}_{{\color[rgb]{0.453125,0,0}+8.76}}$
-XCLIP-FT AP 93.32 100.00 99.97 85.55 98.97 99.60 99.98 100.00 97.77 95.75 - $\textbf{97.10}_{{\color[rgb]{0.453125,0,0}+10.43}}$
@@
Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Real
@@
Pika NPR Image ACC 55.36 77.57 71.88 4.86 7.21 4.29 86.26 60.29 71.43 31.53 99.52 51.83
AP 45.74 91.55 92.71 21.80 44.32 22.74 95.04 90.03 84.91 60.88 - 64.97
STIL Image ACC 75.00 79.43 94.49 57.86 53.14 64.14 97.12 85.29 69.43 62.42 92.43 75.52
AP 22.35 71.62 93.19 40.61 53.24 47.73 94.94 85.82 58.99 61.91 - 63.04
XCLIP-B-FT Video ACC 67.86 91.29 96.23 12.00 22.36 9.14 99.84 83.43 75.57 51.84 99.64 64.47
AP 71.08 97.53 99.44 44.68 72.69 38.37 99.96 97.32 88.00 74.00 - 78.31
DeMamba Video ACC 92.86 97.29 98.48 38.29 53.50 41.43 99.84 94.07 77.29 64.15 98.65 $\textbf{77.80}_{{\color[rgb]{0.453125,0,0}+13.33}}$
-XCLIP-FT AP 77.75 98.42 99.16 52.97 76.72 56.24 99.80 97.91 82.83 74.81 - $\textbf{81.66}_{{\color[rgb]{0.453125,0,0}+3.35}}$
SEINE NPR Image ACC 46.43 78.57 63.70 21.86 7.00 3.29 92.97 89.29 33.86 24.84 99.70 51.05
AP 36.30 92.63 85.02 52.68 25.69 11.05 97.80 97.78 64.64 47.48 - 61.11
STIL Video ACC 71.43 80.43 88.48 67.71 54.57 55.71 93.93 89.57 72.00 50.11 92.27 74.20
AP 23.89 71.01 88.18 52.17 54.49 41.23 84.73 87.38 58.72 46.51 - 60.83
XCLIP-B-FT Video ACC 85.71 95.43 76.23 65.86 35.93 37.00 99.68 99.00 75.57 49.78 99.80 74.54
AP 85.89 97.97 94.40 92.81 81.68 77.68 98.48 98.91 92.27 67.91 - 88.80
DeMamba Video ACC 94.64 98.43 92.17 82.43 52.29 54.00 99.52 99.14 79.29 57.88 98.99 $\textbf{82.61}_{{\color[rgb]{0.453125,0,0}+8.07}}$
-XCLIP-FT AP 83.74 99.01 97.66 90.82 84.11 73.30 99.72 99.73 89.72 76.45 - $\textbf{89.43}_{{\color[rgb]{0.453125,0,0}+0.63}}$
OpenSora NPR Image ACC 55.36 76.29 55.51 58.57 76.50 22.43 74.92 83.07 29.86 60.37 95.95 62.62
AP 25.65 75.24 65.02 55.12 82.42 20.75 72.65 86.84 28.13 64.50 - 57.63
STIL Video ACC 32.14 45.43 56.45 35.14 45.07 34.57 57.83 63.14 19.86 43.95 98.13 48.33
AP 6.94 55.62 75.99 43.55 68.06 44.01 63.84 80.59 29.39 57.58 - 52.56
XCLIP-B-FT Video ACC 67.86 75.86 67.46 70.86 73.14 43.57 79.87 86.29 33.43 63.17 98.10 69.06
AP 48.28 81.39 81.84 77.38 86.08 51.87 83.41 93.18 39.27 72.74 - 71.54
DeMamba Video ACC 55.36 87.43 81.30 73.14 85.21 73.14 89.62 90.07 44.86 58.10 97.30 $\textbf{75.95}_{{\color[rgb]{0.453125,0,0}+6.89}}$
-XCLIP-FT AP 25.89 86.63 87.27 74.38 91.12 76.01 86.41 93.83 48.74 67.92 - $\textbf{73.82}_{{\color[rgb]{0.453125,0,0}+2.28}}$

### v3

@@ 3.3 Video collection details of GenVideo
Fake video collection: The guiding principle for collecting fake videos is to ensure maximal diversity in content and generators. We prioritize generating additional fake videos using the most recent generators due to their superior quality. To collect diverse fake videos from different resources as training samples, we have established a video generation pipeline for text-to-video generation and image-to-video generation. This pipeline facilitates the production of videos using popular generative mechanisms, including diffusion-based models (Xing et al., 2023; Zhang et al., 2023b; Guo et al., 2023a; Wang et al., 2023c; Blattmann et al., 2023b; Chen et al., 2024; Xing et al., 2023; Zhang et al., 2023b; Chen et al., 2023c), transformer-based methods (Ma et al., 2024b; Ope, 2024) and service-based method Pika website (pik, 2022). For image-to-video generation, we employed various text-to-image models to produce diverse images, including different versions of Stable Diffusion (SD (Rombach et al., 2022), SDXL (Podell et al., 2023)). In order to produce videos with rich semantics through different generative approaches, including semantic-diverse text and image prompts across person, objects and diverse scenes, we firstly construct a rich prompt dictionary. In detail, we selected common 100 categories such as humans, animals, plants and etc as foreground keywords, and typical 20 scenes like “in the park” or “on the lawn” as background keywords. Leveraging a large language model (Le Scao et al., 2023), these foreground and background keywords are expanded into about 4000 comprehensive textual prompts. Besides, we also randomly sample 1600 textual prompts from VBench Huang et al. (2024) with consideration of semantic diversity and style diversity. Secondly, each prompt from constructed prompt dictionary is fed into T2V (resp., T2I) generators to produce diverse videos (resp., images). I2V generators leverage images generated from T2I generators to produce videos.
@@ 5.1 Implementation details
Model Batchsize (Video level) LR Epochs Optimizer Scheduler Loss
F3Net 128 1e-5 30 AdamW [20, 25], lr * 0.1 BCE
NPR 128 1e-5 30 [20, 25], lr * 0.1
STIL 128 1e-5 30 [20, 25], lr * 0.1
VideoMAE-B 128 1e-5 30 [20, 25], lr * 0.1
CLIP-B-LP 128 1e-6 30 [20, 25], lr * 0.1
DeMamba-CLIP-LP 128 1e-6 30 [20, 25], lr * 0.1
XCLIP-B-LP 128 1e-6 30 [20, 25], lr * 0.1
DeMamba-XCLIP-LP 128 1e-6 30 [20, 25], lr * 0.1
CLIP-B-FT 128 1e-6 10 -
MINTIME-CLIP-B 128 1e-6 10 -
FTCN-CLIP-B 128 1e-6 10 -
TALL 128 1e-6 10 -
XCLIP-B-FT 128 1e-6 10 -
DeMamba-CLIP-FT 128 1e-6 10 -
DeMamba-XCLIP-FT 128 1e-6 10 -
Table 3: Comparisons to the SOTAs in F1 score (F1) and average precision (AP) on the many-to-many generalization task.
Model Detection Metric Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild Avg.
@@
F3Net Image R 0.8393 0.9971 0.9862 0.7757 0.5700 0.3657 0.9952 0.9971 0.8943 0.7678 0.8188
F1 0.5000 0.9406 0.9628 0.8169 0.6988 0.4904 0.9332 0.9688 0.8873 0.8251 0.8024
AP 0.6827 0.9989 0.9967 0.8935 0.8524 0.6317 0.9958 0.9989 0.9380 0.8841 0.8873
NPR Image R 0.9107 0.9957 0.9949 0.2429 0.8964 0.5771 0.9712 0.9986 0.9429 0.8780 0.8408
F1 0.2786 0.8441 0.9131 0.3028 0.8627 0.5944 0.8170 0.9164 0.8184 0.8163 0.7164
AP 0.6717 0.9914 0.9920 0.2276 0.9391 0.6176 0.9633 0.9972 0.9415 0.9040 0.8245
STIL Video R 0.7857 0.9814 0.9804 0.7600 0.6179 0.5329 0.9936 0.9736 0.9457 0.6501 0.8222
F1 0.3805 0.9068 0.9458 0.7824 0.7232 0.6217 0.9039 0.9433 0.8884 0.7267 0.7823
AP 0.5721 0.9908 0.9932 0.8619 0.8224 0.7043 0.9925 0.9896 0.9718 0.8132 0.8712
VideoMAE-B Video R 0.6786 0.9600 0.9841 0.9614 0.7714 0.8043 0.9744 0.9693 0.9629 0.6836 0.8750
F1 0.623 0.9593 0.9819 0.9600 0.8608 0.8722 0.9644 0.9745 0.9615 0.7977 0.8955
AP 0.6649 0.9885 0.9977 0.9927 0.9655 0.9531 0.9949 0.9969 0.9927 0.9074 0.9454
MINTIME-CLIP-B Video R 0.8929 1.0000 0.9899 0.2643 0.9679 0.9814 0.9984 1.0000 0.8429 0.8238 0.8762
F1 0.4902 0.9340 0.9606 0.3760 0.9499 0.9246 0.9266 0.9659 0.8495 0.8539 0.8231
AP 0.8321 0.9999 0.9967 0.5084 0.9920 0.9927 0.9976 0.9999 0.9183 0.9177 0.9155
FTCN-CLIP-B Video R 0.8750 1.0000 0.9891 0.1771 0.9771 0.9186 1.0000 1.0000 0.8529 0.8283 0.8618
F1 0.7840 0.9859 0.9873 0.2922 0.9813 0.9442 0.9843 0.9929 0.9073 0.8960 0.8755
AP 0.9179 0.9999 0.9979 0.4594 0.9976 0.9780 0.9999 0.9999 0.9469 0.9232 0.9221
TALL Video R 0.9107 0.9828 0.9783 0.8300 0.7657 0.7957 0.9952 0.9893 0.9414 0.6631 0.8852
F1 0.2615 0.8240 0.8964 0.7430 0.7782 0.7234 0.8133 0.9032 0.8027 0.6736 0.7419
AP 0.7115 0.9689 0.9851 0.7938 0.8459 0.7938 0.9879 0.9902 0.9270 0.7647 0.8791
CLIP-B-PT Image R 0.8571 0.8243 0.9036 0.7100 0.7929 0.7543 0.8962 0.8629 0.8214 0.7516 0.8174
F1 0.0215 0.2078 0.1816 0.3271 0.1918 0.2053 0.3509 0.2068 0.2357 0.2289 0.2157
AP 0.6780 0.4356 0.7088 0.2997 0.5297 0.3536 0.5552 0.6603 0.4423 0.4299 0.4483
DeMamba-CLIP-PT Video R 0.5893 0.9643 0.9312 0.6800 0.6936 0.6900 0.8914 0.9186 0.9614 0.5659 $0.7886_{{\color[rgb]{0,0.453125,0}-0.0308}}$
F1 0.234 0.8610 0.8992 0.6949 0.7574 0.7015 0.8105 0.8934 0.8588 0.6379 $0.7349_{{\color[rgb]{0.453125,0,0}+0.5060}}$
AP 0.2587 0.9514 0.9623 0.7343 0.8331 0.7549 0.9017 0.9506 0.9505 0.6995 $0.7997_{{\color[rgb]{0.453125,0,0}+0.3514}}$
CLIP-B-FT Image R 0.9464 0.9986 0.9138 0.7729 0.8814 0.8600 0.9968 0.9979 0.8429 0.8467 0.9057
F1 0.2818 0.8422 0.8691 0.7204 0.8521 0.7698 0.8252 0.9137 0.7608 0.7955 0.7631
AP 0.8067 0.9967 0.9524 0.8220 0.9348 0.8862 0.9955 0.9979 0.8693 0.8908 0.9152
DeMamba-CLIP-B-FT Video R 0.9571 1.0000 0.9870 0.6914 0.9243 0.9329 1.0000 1.0000 0.8357 0.8294 $0.9158_{{\color[rgb]{0.453125,0,0}+0.0101}}$
F1 0.6463 0.9615 0.9739 0.7803 0.9414 0.9276 0.9572 0.9804 0.8723 0.8782 $0.8919_{{\color[rgb]{0.453125,0,0}+0.1288}}$
AP 0.8550 1.0000 0.9959 0.7615 0.9678 0.9699 0.9997 1.0000 0.8980 0.8972 $0.9345_{{\color[rgb]{0.453125,0,0}+0.1930}}$
XCLIP-B-PT Video R 0.8134 0.8215 0.8335 0.8098 0.8182 0.8155 0.8214 0.8298 0.8193 0.8110 0.8193
F1 0.0458 0.3666 0.5237 0.3619 0.5196 0.3640 0.3424 0.5250 0.3656 0.4240 0.3839
AP 0.1639 0.7216 0.8777 0.3986 0.6557 0.5426 0.7523 0.8480 0.6160 0.5528 0.6129
DeMamba-XCLIP-PT Video R 0.6607 0.9586 0.9464 0.7786 0.7536 0.8029 0.9089 0.9250 0.9600 0.6641 $0.8359_{{\color[rgb]{0.453125,0,0}+0.0166}}$
F1 0.1244 0.7223 0.8232 0.6293 0.7172 0.6427 0.6758 0.8137 0.7226 0.6058 $0.6477_{{\color[rgb]{0.453125,0,0}+0.2638}}$
AP 0.1826 0.9350 0.9472 0.6994 0.7808 0.7150 0.8395 0.9223 0.9354 0.6810 $0.7638_{{\color[rgb]{0.453125,0,0}+0.1509}}$
XCLIP-B-FT Video R 0.8214 0.9957 0.9362 0.6129 0.7936 0.6971 0.9792 0.9979 0.7714 0.8359 0.8441
F1 0.3147 0.8805 0.9041 0.6530 0.8242 0.7099 0.8602 0.9370 0.7570 0.8212 0.7662
AP 0.6442 0.9973 0.9678 0.7098 0.9035 0.7728 0.9734 0.9984 0.8201 0.8897 0.8677
DeMamba-XCLIP-FT Video R 0.9821 1.0000 0.9986 0.6543 0.9486 0.9886 1.0000 1.0000 0.9286 0.8909 $\textbf{0.9392}_{{\color[rgb]{0.453125,0,0}+0.0951}}$
F1 0.6467 0.9602 0.9790 0.7539 0.9537 0.9551 0.9557 0.9797 0.9240 0.9120 $\textbf{0.9020}_{{\color[rgb]{0.453125,0,0}+0.1358}}$
AP 0.9332 1.0000 0.9997 0.8555 0.9897 0.9960 0.9998 1.0000 0.9777 0.9575 $\textbf{0.9710}_{{\color[rgb]{0.453125,0,0}+0.1043}}$
@@
Sora Morph Gen2 HotShot Lavie Show-1 Moon Crafter Model Wild
@@
Pika NPR Image R 0.5536 0.7757 0.7188 0.4860 0.0721 0.0429 0.8626 0.6029 0.7143 0.3153 0.5144
F1 0.4627 0.8410 0.8197 0.0871 0.1293 0.0772 0.8894 0.7368 0.8019 0.4604 0.5306
AP 0.4574 0.9155 0.9271 0.2180 0.4432 0.2274 0.9504 0.9003 0.8491 0.6088 0.6497
STIL Image R 0.7500 0.7943 0.9449 0.5786 0.5314 0.6414 0.9712 0.8529 0.6943 0.6242 0.7383
F1 0.0984 0.5530 0.7580 0.4355 0.5128 0.4708 0.6107 0.7130 0.5008 0.5117 0.5165
AP 0.2235 0.7162 0.9319 0.4061 0.5324 0.4773 0.9494 0.8582 0.5899 0.6191 0.6304
MINTIME-CLIP-B Video R 0.1607 0.2686 0.1471 0.0557 0.3757 0.0900 0.2188 0.4071 0.0114 0.4309 0.2166
F1 0.2338 0.4178 0.2545 0.1039 0.5428 0.1626 0.3535 0.5752 0.0222 0.5969 0.3263
AP 0.1692 0.7433 0.6066 0.3903 0.8229 0.4108 0.7479 0.8862 0.2385 0.7225 0.5738
FTCN-CLIP-B Video R 0.9107 0.9300 0.9232 0.1771 0.1750 0.0643 0.9968 0.8479 0.7457 0.5173 0.6287
F1 0.5896 0.9195 0.9371 0.2770 0.2855 0.1088 0.9483 0.8948 0.8106 0.6513 0.6423
AP 0.8369 0.9705 0.9718 0.3747 0.4490 0.1733 0.9975 0.9541 0.7971 0.6686 0.7194
TALL Video R 0.6071 0.7043 0.9449 0.5914 0.5757 0.6400 0.9633 0.8871 0.6057 0.6210 0.7141
F1 0.1038 0.5609 0.8029 0.4932 0.5817 0.5231 0.6726 0.7746 0.5024 0.5569 0.5572
AP 0.1582 0.6363 0.9347 0.4400 0.5907 0.5106 0.9209 0.8785 0.5111 0.6443 0.6225
XCLIP-B-FT Video R 0.6786 0.9129 0.9623 0.1200 0.2236 0.0914 0.9984 0.8343 0.7557 0.5184 0.6096
F1 0.5891 0.9301 0.9679 0.2049 0.3581 0.1579 0.9712 0.8974 0.8361 0.6662 0.6579
AP 0.7108 0.9753 0.9944 0.4468 0.7269 0.3837 0.9996 0.9732 0.8800 0.7400 0.7831
DeMamba-XCLIP-FT Video R 0.9286 0.9729 0.9848 0.3829 0.5350 0.4143 0.9984 0.9407 0.7729 0.6415 $\textbf{0.7572}_{{\color[rgb]{0.453125,0,0}+0.1476}}$
F1 0.4298 0.8990 0.9460 0.4864 0.6559 0.5160 0.9017 0.9235 0.7869 0.7183 $\textbf{0.7263}_{{\color[rgb]{0.453125,0,0}+0.0684}}$
AP 0.7775 0.9842 0.9916 0.5297 0.7672 0.5624 0.9980 0.9791 0.8283 0.7481 $\textbf{0.8166}_{{\color[rgb]{0.453125,0,0}+0.0335}}$
SEINE NPR Image R 0.4643 0.7857 0.6370 0.2186 0.0700 0.0329 0.9297 0.8929 0.3386 0.2484 0.4618
F1 0.4685 0.8592 0.7684 0.3469 0.1283 0.0612 0.9401 0.9332 0.4907 0.3882 0.5385
AP 0.3630 0.9263 0.8502 0.5268 0.2569 0.1105 0.9780 0.9778 0.6464 0.4748 0.6111
STIL Video R 0.7143 0.8043 0.8848 0.6771 0.5457 0.5571 0.9393 0.8957 0.7200 0.5011 0.7239
F1 0.0922 0.5533 0.7240 0.4864 0.5199 0.4181 0.5921 0.7317 0.5099 0.4292 0.5057
AP 0.2389 0.7101 0.8818 0.5217 0.5449 0.4123 0.8473 0.8738 0.5872 0.4651 0.6083
MINTIME-CLIP-B Video R 0.6964 0.9257 0.9116 0.1071 0.2293 0.1071 0.9968 0.8600 0.7371 0.5637 0.6135
F1 0.5532 0.9297 0.9378 0.1827 0.3633 0.1827 0.9630 0.9087 0.8177 0.6988 0.6538
AP 0.6596 0.9781 0.9790 0.3688 0.6215 0.3311 0.9978 0.9701 0.7963 0.7402 0.7443
FTCN-CLIP-B Video R 0.8929 0.9786 0.6297 0.4043 0.1650 0.1557 0.9984 0.9971 0.6514 0.4060 0.6279
F1 0.8130 0.9765 0.7664 0.5660 0.2802 0.2636 0.9858 0.9922 0.7775 0.5697 0.6991
AP 0.8836 0.9972 0.8736 0.8028 0.6072 0.4321 0.9998 0.9998 0.8139 0.6255 0.8036
TALL Video R 0.5714 0.7643 0.7957 0.6443 0.3486 0.5629 0.8514 0.9143 0.6714 0.4438 0.6568
F1 0.1730 0.7053 0.7957 0.6289 0.4498 0.5727 0.7403 0.86423 0.6469 0.5077 0.6085
AP 0.2716 0.7836 0.8838 0.6896 0.5518 0.6093 0.8240 0.9426 0.7007 0.5483 0.6805
XCLIP-B-FT Video R 0.8571 0.9543 0.7623 0.6586 0.3593 0.3700 0.9968 0.9900 0.7557 0.4978 0.7201
F1 0.7705 0.9632 0.8580 0.7814 0.5234 0.5291 0.9826 0.9879 0.8468 0.6548 0.7898
AP 0.8589 0.9797 0.9440 0.9281 0.8168 0.7768 0.9848 0.9891 0.9227 0.6791 0.8880
DeMamba-XCLIP-FT Video R 0.9464 0.9843 0.9217 0.8243 0.5229 0.5400 0.9952 0.9914 0.7929 0.5788 $\textbf{0.8098}_{{\color[rgb]{0.453125,0,0}+0.0897}}$
F1 0.5000 0.9255 0.9240 0.8381 0.6559 0.6412 0.9228 0.9609 0.8192 0.6855 $\textbf{0.7873}_{{\color[rgb]{0,0.453125,0}-0.0025}}$
AP 0.8374 0.9901 0.9766 0.9082 0.8411 0.7330 0.9972 0.9973 0.8972 0.7645 $\textbf{0.8943}_{{\color[rgb]{0.453125,0,0}+0.0063}}$
OpenSora NPR Image R 0.5536 0.7629 0.5551 0.5857 0.7650 0.2243 0.7492 0.8307 0.2986 0.6037 0.5929
F1 0.1263 0.6520 0.6008 0.5406 0.7448 0.2490 0.6248 0.7835 0.3184 0.5918 0.5232
AP 0.2565 0.7524 0.6502 0.5512 0.8242 0.2075 0.7265 0.8684 0.2813 0.6450 0.5763
STIL Video R 0.3214 0.4543 0.5645 0.3514 0.4507 0.3457 0.5783 0.6314 0.1986 0.4395 0.4336
F1 0.1313 0.5282 0.6644 0.4332 0.5686 0.4277 0.6167 0.7153 0.2712 0.5349 0.4892
AP 0.0694 0.5562 0.7599 0.4355 0.6806 0.4401 0.6384 0.8059 0.2939 0.5758 0.5256
MINTIME-CLIP-B Video R 0.8571 0.9214 0.7971 0.6343 0.2357 0.3057 0.9712 0.9636 0.7257 0.4773 0.6889
F1 0.6000 0.9208 0.8675 0.7396 0.3695 0.4412 0.9426 0.9619 0.8044 0.6212 0.7269
AP 0.7981 0.9739 0.9370 0.8451 0.6232 0.5819 0.9891 0.9926 0.8529 0.6925 0.8286
FTCN-CLIP-B Video R 0.1964 0.1786 0.1442 0.1957 0.1579 0.0329 0.2716 0.3229 0.0214 0.3413 0.1863
F1 0.3235 0.3027 0.2519 0.3270 0.2725 0.0635 0.4252 0.4879 0.0419 0.5084 0.3005
AP 0.3178 0.6675 0.4927 0.6954 0.7089 0.2382 0.7876 0.8500 0.2263 0.7049 0.5689
TALL Video R 0.3393 0.5771 0.5225 0.3986 0.5893 0.6129 0.5655 0.6736 0.2429 0.3952 0.4916
F1 0.1357 0.6173 0.6256 0.4713 0.6790 0.6432 0.5975 0.7405 0.3163 0.4890 0.5315
AP 0.0798 0.6447 0.7246 0.4924 0.7830 0.6916 0.6166 0.8329 0.3132 0.5313 0.5710
XCLIP-B-FT Video R 0.6786 0.7586 0.6746 0.7086 0.7314 0.4357 0.7987 0.8629 0.3343 0.6317 0.6615
F1 0.2686 0.7479 0.7443 0.7162 0.7833 0.5096 0.7595 0.8638 0.4167 0.6875 0.6497
AP 0.4828 0.8139 0.8184 0.7738 0.8608 0.5187 0.8341 0.9318 0.3927 0.7274 0.7154
DeMamba-XCLIP-FT Video R 0.5536 0.8743 0.8130 0.7314 0.8521 0.7314 0.8962 0.9007 0.4486 0.5810 $\textbf{0.7382}_{{\color[rgb]{0.453125,0,0}+0.0767}}$
F1 0.1742 0.7742 0.8094 0.6905 0.8333 0.6905 0.7706 0.8604 0.4895 0.6209 $\textbf{0.6713}_{{\color[rgb]{0.453125,0,0}+0.0216}}$
AP 0.2589 0.8663 0.8727 0.7438 0.9112 0.7601 0.8641 0.9383 0.4874 0.6792 $\textbf{0.7382}_{{\color[rgb]{0.453125,0,0}+0.0228}}$